### v1

# Hardness-Dependent Optimized Quantum Adiabatic Schedules for Maximum Independent Sets

Abstract
We propose a numerical approach to design highly efficient schedules for adiabatic quantum computing, focusing on the maximum independent set problem and on neutral atom platforms. Based on a representative dataset of small graphs, we present numerical evidences that the optimum schedules depend principally on the hardness of the problem and not on its size. These schedules perform better than the benchmark protocols and admit a straightforward implementation in the hardware. This allows us to extrapolate the results to larger graphs and to successfully solve moderately hard instances with more than 100 atoms using QuEra’s quantum computer “Aquila”. We believe that extending our approach to hybrid algorithms could be the key to solve the hardest problems with the current technology.
maximum independent set; neutral atoms; QuEra’s Aquila; analog quantum computing
I Introduction
In recent years, quantum computers have emerged as a promising way to solve complex optimization tasks in science and industry [1, 2, 3, 4]. Typically, the problem of interest is encoded in the ground state of a quantum Hamiltonian that can be reached through adiabatic evolution, starting from an easy-to-prepare initial state [5, 6, 7, 8]. Advanced protocols have to be developed to overcome not only the experimental limitations, such as qubit decoherence and control errors, but also the intrinsic complication of vanishing gap in the eigenspectrum during the evolution [9, 10, 11, 12].
Here, we focus on the maximum independent set (MIS) problem, where the task is to find the largest set of vertices that do not share an edge in a graph. Such problems appear naturally in practical applications, for instance while designing resilient communication networks [13], and are known to be NP-hard in general [14]. Finding a MIS is still a difficult task in the particular case of unit-disk graphs, in which vertices are connected by an edge if they are closer than a fixed distance [15]. This family of graphs admits a direct mapping to arrays of Rydberg atoms, which explains the popularity of the MIS problem in analog quantum computing [16, 17, 18].
In this setting, optical tweezers enable arbitrary arrangements of atoms and controlled laser pulses manipulate the quantum states [19, 20]. These lasers drive the adiabatic evolution of the system following schedules that are specific to the MIS problem. The precise description of the schedules is of utmost importance since it conditions the success of the quantum algorithm. Many methods have been proposed in that regard: quantum approximate optimization algorithm and variational quantum eigensolver [21, 22, 23, 24], shortcut to adiabadicity and counterdiabatic protocol [25, 26, 27], Bayesian optimization and machine learning [28, 29], to name a few.
Most aforementioned methods require multiple iterations to be run on a quantum device and to be updated by a classical optimizer and thus belong to the class of hybrid algorithms [30]. By their very nature, these algorithms cost both time and money as they make intense use of the quantum hardware. In this study, we show how classical precomputation of optimized schedules for small graphs can lead to highly efficient protocols for large graphs on the quantum device. First, we introduce the mathematical concepts that are used throughout this work. Then, in Sec. II we describe how to optimize the schedules based on numerical simulations of small graphs, and in Sec. III we present the results obtained in experiments. Detailed calculations and additional results are presented in the appendix.
Rydberg Hamiltonian
The analog quantum dynamics is governed by the Hamiltonian $H=H_{\text{drive}}+H_{\text{cost}}$, where the driving part and the cost function are given by [18]
[…]
$\frac{H_{\text{cost}}(t)}{\hbar}$ $=-\Delta(t)\sum_{i}\hat{n}_{i}+\sum_{i<j}V_{ij}\hat{n}_{i}\hat{n}_{j}.$ (1b)
In these equations, $\Omega$ is the Rabi drive amplitude, $\phi$ its phase, $\Delta$ the detuning, and $\hat{n}_{i}=\ket{1}_{i}\bra{1}$ | detects the Rydberg excitation of the atom $i=1,\ldots,N$. The two-body potential is $V_{ij}=C_{6}/|\textbf{r}_{i}-\textbf{r}_{j}|^{6}$, with $C_{6}$ a constant for the Van der Waals interaction between two Rydberg states at positions $\textbf{r}_{i}$ and $\textbf{r}_{j}$, which encodes the unit-disk graph of the MIS problem to be solved. The boundary conditions at initial and final times $t_{\text{i}}$ and $t_{\text{f}}$ for an adiabatic evolution are usually stated as follows:
$\Omega_{\text{i}}=\Omega_{\text{f}}=0,$ (2a)
$\Delta_{\text{i}}<0<\Delta_{\text{f}}.$ (2b)
In practice, we implement the first inequality as $\Delta_{\text{i}}\leq-\delta_{\Delta}$, where $\delta_{\Delta}$ is a small positive constant that incorporates the main sources of error in the quantum hardware, so that all atoms are ensured to start in the ground state |$\ket{0}$. More importantly, we show in Sec. SI that the final detuning should lie in a narrow interval $[\Delta_{\text{LB}},\Delta_{\text{UB}}]$, where the bounds are explicitly given by
$\frac{C_{6}}{12\,a^{6}}\leq\Delta_{\text{f}}\leq\frac{C_{6}}{8\,a^{6}},$ (3)
with $a$ the lattice spacing. Note that these bounds are not universal, but they are satisfied by the vast majority of unit-disk graphs on a square lattice.
Quantum hardware
[…]
Hardness parameter
A detailed analysis of the dynamics of the (classical) simulated annealing for MIS reveals that the graphs with many suboptimal solutions are likely to trap the algorithm in local minima, hence missing the global solution [18]. This observation leads to the following definition of the hardness parameter of a graph:
$\mathcal{HP}=\frac{D_{|\text{MIS}|-1}}{|\text{MIS}|\cdot D_{|\text{MIS}|}},$ (4)
[…]
II Numerical optimization
In this section, we consider a fixed set of 500 representative graphs of size 8 to 17  [^1] We equally refer to the number of vertices in a graph as order or size. In quantum computing, the latter is commonly understood as the size of the system and not as the number of edges in the graph (mathematical definition)., which allows us to perform fast classical simulations and to use efficient deterministic optimization algorithms. This dataset is generated in such a way that the correlation between the size of the graphs and their hardness parameter is minimized, see Sec. SIII. Therefore, the results that depend principally on $\mathcal{HP}$ may be extrapolated, in principle, to graphs of any size. We focus on the probability to reach a MIS solution at the end of the adiabatic evolution, which we denote by $P_{\text{MIS}}$ and use as the score function to be optimized. Note that most studies consider the approximation ratio instead (normalized mean energy of the final state), but in our opinion only the lowest energy level and not the full energy spectrum has a practical utility because of the very definition of the hardness parameter, see Eq. (4). More details on the dataset and the numerical methods are given in Sec. SIII.
Following the “best practices” listed in [20], we consider protocols for which the Rabi drive $\Omega(t)$ reaches $\Omega_{\max}$ during the evolution, so that the associated dynamic blockade radius becomes $R_{\text{b}}=(C_{6}/\Omega_{\max})^{1/6}\approx 8.367$. In order to generate a unit-disk graph on nearest and next-nearest neighbors, this radius should satisfy $\sqrt{2}a<R_{\text{b}}<2a$, which in turn implies $5.9164.183<a<$ 5.916. We set $a=5$ which yields good results in preliminary calculations. We also consider a fixed evolution time $t_{\text{f}}=1$ so that $P_{\text{MIS}}$ varies as much as possible depending on the various protocols; much smaller or larger times would lead to either extremely bad or extremely good results independently of the precise shape of the schedules.
In the following sections, we optimize and compare two paradigmatic protocols based on either piecewise linear or counterdiabatic schedules.
II.1 Piecewise linear protocol
The typical adiabatic protocol is based on piecewise linear schedules, where $\Omega$ is slowly ramped up, then $\Delta$ is ramped from negative to positive, and finally $\Omega$ is ramped down, see [20, Fig. 3.1b] or [34, Fig. 2c]. Usually, these schedules are symmetric in the sense that the initial and final ramp times $\tau_{\text{i}}$ and $\tau_{\text{f}}$ are equal, and that the initial and final detunings $\Delta_{\text{i}}$ and $\Delta_{\text{f}}$ have opposite values. In our case, we are looking for schedules that are as general as possible, so we keep all parameters independent. Moreover, we consider distinct values of the ramp times for $\Omega$ and $\Delta$. After few optimization steps, it appears that the ramp times of the detuning tend to vanish; we therefore set them to zero. Then, we optimize the four remaining parameters $(\tau_{\text{i}},\tau_{\text{f}},\Delta_{\text{i}},\Delta_{\text{f}})$ for each graph of the dataset. One interesting result is that the optimum parameters mostly depend on the hardness parameter and not on the order of the graph, see Fig. 1. The consequence is twofold: first, we can fit these data to create a simple model of optimized schedules, se Fig. 2, and second, the results obtained from small graphs can be extrapolated to larger graphs (Sec. III.2). In other words, many costly and time consuming calls to quantum computers that are required in hybrid optimization may be saved by using parameters that are designed by purely classical means.
Figure 1: Optimum parameters of the linear schedules for the 500 representative graphs. These values depend on $\mathcal{HP}$ and converge quickly to a constant for sufficiently hard graphs. This behavior is particularly strong for $\tau_{\text{f}}$ (top) and $\Delta_{\text{f}}$ (bottom) at the end of the adiabatic evolution. There is no statistical difference between the smallest graphs (down purple triangles) and the largest ones (up yellow-green triangles).
Figure 2: Optimized piecewise linear schedules. These functions depend on the hardness of the graph and are illustrated here for $\mathcal{HP}\in\{0.5,1,2,4,8\}$ (thin lines, from green to maroon). Any large graph of unknown hardness is best solved by using the parameters of the limiting case $\mathcal{HP}\to\infty$ (dark maroon thick line): $\tau_{\text{i}}\approx t_{\text{f}}/2$, $\tau_{\text{f}}\approx t_{\text{f}}/10$, $\Delta_{\text{i}}\approx-\Delta_{\text{UB}}/2$, $\Delta_{\text{f}}=\Delta_{\text{UB}}$.
A priori knowledge of the hardness parameter
The aim of the adiabatic evolution is to find the MIS, which only then allows one to calculate the value of $\mathcal{HP}$. In the case of small graphs, since all vertex configurations are enumerated easily, we consider that $\mathcal{HP}$ is known and that it can be used to choose the best schedules based on the model. For large graphs, however, the hardness parameter cannot be considered to be a known value. In that case, one should either use the schedules with $\mathcal{HP}\to\infty$, which yields quite good results for all graphs, or eventually consider a one-dimensional search on the hardness parameter, which is still much easier than performing a multi-dimensional optimization of all schedule parameters.
Further improvement
The linear schedules depicted in Fig. 1 yield very good results compared to existing protocols, which motivates us to go one step further by considering not 4 but 6 free parameters, see Sec. SIII.3. The optimized schedules are still easily implementable, yet they perform much better than all benchmark protocols, see Sec. II.3.
II.2 Couterdiabatic protocol
The limited coherence time of the quantum state in the quantum hardware is a major source of error in adiabatic computing. Shortcuts to adiabaticity are well-known methods to overcome such limitations [35]. Among such methods, the counterdiabatic (CD) driving has gained considerable attention over the last years [36, 25, 37]. The idea of the CD driving is to add a gauge potential $\mathcal{A}$ to the Hamiltonian $H$ such that the transitions between eigenstates are suppressed:
$H_{\text{CD}}(t)=H(t)+\mathcal{A}(t).$ (5)
The calculation of the exact gauge potential is a very difficult task in general, and these mathematical terms may not be always adapted to an experimental implementation. Neglecting the interaction terms in the calculations, one can however find a suitable gauge potential which utilizes the phase of the Rabi drive [27]. We generalize the smooth schedules of this so-called Analog Counterdiabatic Quantum Computing (ACQC) protocol as follows:
$\Omega(t)$ $=\Omega_{\max}\,\sin\left(\frac{\pi}{2}\,\sin\left(\theta_{\Omega}(t)\right)\right)^{2},$ (6a)
$\Delta(t)$ $=\frac{\Delta_{\text{i}}+\Delta_{\text{f}}}{2}+\frac{\Delta_{\text{i}}-\Delta_{\text{f}}}{2}\cos\left(\theta_{\Delta}(t)\right),$ (6b)
where $\theta(t)$ is any monotonic smooth function that ranges from 00 to $\pi$; here we set $\theta_{\Omega,\Delta}(t)=\pi\,t/t_{\text{f}}$. Moreover, based on the variational procedure proposed in [26], we show in Sec. SII how to improve the standard ACQC protocol to explicitly take into account the graph-specific coefficients $V_{ij}$. We consider an additional free parameter $0\leq\nu\leq C_{6}/a^{6}$ that interpolates between these two gauge potentials, which greatly helps in the optimization process.
As for the piecewise linear protocol, we optimize the schedule parameters for all graphs of the dataset. We observe the same dependence on the hardness parameter (Fig. 3) as well as a strong shift of the initial and final detunings (Fig. S7). We also tried to improve the results by considering more evolved $\theta$ functions (in particular, to mimic the skewed shape of the Rabi amplitude in Fig. 2), but no significant difference was found in this case.
Hardware implementation
In Aquila, the Rabi amplitude and the detuning must be defined as piecewise linear and not smooth functions, while the phase must be piecewise constant. However, the step size can be as small as 0.05 µs, which results in maximum 80 intervals for a coherent evolution. The schedules are therefore well approximated in the quantum hardware, but one could slightly improve the experimental results by taking this hardware capability into account in the numerical optimization.
Figure 3: Optimum parameters for the CD schedules. The initial detuning (top) reaches the limit $-\delta_{\Delta}$ for several graphs, but in average it tends to $\Delta_{\text{i}}\approx-\Delta_{\text{UB}}/10$ for the hardest graphs. The final detuning (bottom) also increases with the hardness parameter and quickly reaches the limit $\Delta_{\text{UB}}$. Note that the easy graphs exhibit optimized values of $\Delta_{\text{f}}$ that are smaller than the statistically determined constant $\Delta_{\text{LB}}$. This is due to the fact that the exact graph-dependent lower bounds are much smaller in that case, see Fig. S2. Finally, the interpolating coefficient (not shown) displays a similar behavior and tends to a strictly positive value $\nu=4.30$ in the current setting, indicating that the generalized gauge potential plays an important role in this MIS protocol.
II.3 Results
In Fig. 4, we compare the probability $P_{\text{MIS}}$ to reach the solution for 4 protocols: the piecewise linear and ACQC protocols defined in [27] which serve as benchmarks, and their generalized and hardness-dependent counterparts described in the previous sections. Note that $a=5.5$ for the benchmark protocols but $a=5.0$ in our study. The $\mathcal{HP}$-linear protocol yields the best results for all graphs and in particular for the hard ones, with a tenfold improvement over the benchmark protocol in this case. Detailed results for all graphs and protocols are illustrated in Fig. S8.
As for the optimized parameters, we observe that the values of $P_{\text{MIS}}$ depend mostly on the hardness parameter and not on the order of the graphs. For instance, the success probabilities for the $\mathcal{HP}$-linear protocol range from 0.59 ($\mathcal{HP}=4.20$) to 0.99 ($\mathcal{HP}=0.46$) for $N=10$, and from 0.11 ($\mathcal{HP}=9.71$) to 0.98 ($\mathcal{HP}=0.52$) for $N=17$.
Figure 4: Probability to reach a MIS depending on the hardness for the 500 graphs of the dataset. The piecewise linear (blue) and ACQC (orange) protocols serve as benchmarks. For hard graphs, the generalized $\mathcal{HP}$-ACQC (green) and $\mathcal{HP}$-linear (red) protocols developed in this study lead to an average success probability that is improved by approximately 150% and 1 000%, respectively.
III Experiments
Based on thorough numerical simulations, we have shown in the previous section that hardness-dependent schedules lead to high probabilities to reach the MIS solution. Now, we extend these results to the hardware and to graphs that are not part of the dataset used in the optimization. First, we consider a small set of “toy graphs” that we can still simulate classically to ensure that the various protocols are correctly run in Aquila. Then, we apply the models to large graphs (137 vertices) to verify that the hardness parameter and not the order of the graph is indeed the most relevant property when solving the MIS problem.
Measurement shots
For all examples, we use 720 shots (repetitions of the adiabatic evolution) to get enough data for a statistical comparison of the methods. We choose this specific value because it is a multiple of 9, 12, and 16, which are the maximum number of copies of the small graphs ($3\times 3$ and $3\times 4$ lattices with $a=5$ and $a=5.5$) that fit and can be run simultaneously in the hardware. We consider only the shots in which the lattice is correctly filled by atoms. The typical probability of failing to occupy a site in Aquila is $\epsilon_{\text{fill}}=0.007$ [20, Sec. 1.5], and in practice we get $684\pm 7$ and $380\pm 14$ valid shots for the small and the large graphs, respectively.
III.1 Toy graphs
In Fig. 5, we display the probability to reach the MIS solution for the 11 toy graphs defined in [28] and detailed in Sec. SIII.1. As in that article, we set $t_{\text{f}}=0.7$ so that the probabilities of success are distributed around one half. We observe a good agreement between numerical simulations and the corresponding results obtained on hardware, especially for the benchmark protocols. One possible reason for the optimized schedules to yield outcomes that are slightly lower than expected may be the choice of the larger lattice spacing, or the larger gradient in $\Omega$; many other sources of error could explain these results, see [20, Sec. 1.5] for an exhaustive list. Nevertheless, the $\mathcal{HP}$-linear protocol still performs better on the hardware than the linear and CD benchmarks, with a threefold and a twofold mean improvement, respectively.
Figure 5: Probability to reach the MIS solution for the 11 toy graphs. The rectangle bars are numerical results and the error bars are experimental results (95% confidence intervals). The ranking of the 4 studied protocols is the same as in Fig. 4, with $\mathcal{HP}\in[2,3]$ here. The black horizontal line represents the average success probability obtained by the Bayesian optimization in [28, Fig. 9], which performs better than the benchmarks (blue and orange) but not as well as the hardness-dependent protocols (green and red).
III.2 Large graphs
We now turn to large graphs, which are of much higher interest in real applications. We consider three of the four graphs studied in [28, Fig. 10], which consist of 137 vertices and whose hardness parameters span three orders of magnitude ($\mathcal{HP}\approx$ 14, 125, and 1435, with the second one referred as “95” in the mentioned article). In addition, we generate one very easy graph of the same size to span yet another order of magnitude ($\mathcal{HP}\approx$ 1.4), see Fig. 6. The hardness parameter of these large graphs is efficiently computed by advanced classical algorithms, namely by the GenericTensorNetworks Julia library [38, 39].
Since the MIS solutions and consequently the hardness parameters are not known a priori, we apply our optimized linear schedules in the limit $\mathcal{HP}\to\infty$. Moreover, we set $t_{\text{f}}=t_{\text{max}}=4$ so that the probability to find a MIS solution is as high as possible. Finally, we post-process all experimental data such that the measured outcomes become valid maximal independent sets. To this end, we use the greedy procedure described in [18, Sec. S2.3]: first, we remove any of two neighboring vertices until the sets become independent. Second, we add vertices to the free sites with no neighbor until the sets are maximal. The same procedure applied to an empty configuration is used as a classical benchmark, which highlights the relative gain obtained with the quantum protocols [20, Sec. 6].
Results are summarized in Fig. 6: the MIS solution is never found by the classical greedy algorithm and is obtained by the standard linear protocol only for the easiest graph. On the contrary, the $\mathcal{HP}$-linear protocol leads to the solution for all but the hardest graph, which indicates a clear improvement over the existing protocols.
IV Discussion and outlook
In this work, we have shown how to carefully design the adiabatic schedules of the Rydberg Hamiltonian to efficiently solve the MIS problem on unit-disk graphs.
First, a statistical investigation of the conditions for which the final quantum state encodes the MIS solution has revealed that the final detuning should lie in a narrow interval. To our knowledge, these are the tightest bounds ever mentioned. We believe that choosing a final detuning in that interval would improve the results of many other studies, in terms of success probability or number of optimization steps. In fact, applying adiabatic protocols in which the final ground state does not correspond to the sought solution can only lead to poorly performing algorithms.
Second, the individual optimization of numerous small graphs has shown that the optimum parameters of the schedules depend mostly on the hardness parameter and only marginally on the size of the graph. Through extensive classical simulations of small systems, we have generated adiabatic schedules that perform extremely well also when applied to much larger graphs; only the hardest instances are not solved by our protocol. Obviously, improvements in the hardware would greatly help in this regard, especially if it permits longer coherent evolution times. Since the quantum technology evolves rapidly, this could be the case soon. However, because of the combinatorial nature of the hardness parameter, very hard MIS problems exist only in sufficiently large graphs, so that our numerical study should be extended to larger graphs to address this issue. We believe that improved schedules could be easily found by considering slightly larger systems, for instance up to about 25 vertices, or even larger using tensor network techniques whenever appropriate [40].
We have focused our work on two families of schedules, namely piecewise linear and smooth counterdiabatic functions. The phase of the Rabi drive is non constant only in the latter case, and we have developed a generalized version of the ACQC protocol [27] to make it specific to each graph. However, when correctly parametrized, the simpler $\mathcal{HP}$-linear protocol performs better in almost all numerical simulations. The difference is even more pronounced in experiments, as the schedules have to be described in terms of piecewise linear or constant functions. We have used a maximum number of 6 free parameters in the optimization, but more general functions may be investigated, as the method described in this work applies to any kind of schedules. Moreover, it would be interesting to develop models of optimized schedules that do not depend only on the hardness parameter but also on other global properties of the graphs, e.g., the mean degree or the mean eccentricity of the vertices. Based on machine learning, such additional predictors could in fact improve the fitting of the schedules.
We think that hybrid classical-quantum algorithms could take advantage of our protocols and, more generally, of precomputed optimized schedules as described in this study: instead of searching in the huge space of all parameters, one could consider lower-dimensional refined ansatzes, such as the $\mathcal{HP}$-parametrization of the schedules or the principal component of the residuals of the fitted parameters.
Finally, we believe that the piecewise linear protocols developed in this study should serve as a new benchmark for any future adiabatic algorithm that aims to solve the MIS problem with Rydberg atoms. In particular, the most basic version of this protocol with 4 free parameters and in the limit of infinite hardness admits a very simple mathematical description, see Fig. 2, and it leads to relatively good results regardless of the specific graph to be solved.
The supporting data for this article are openly available on GitHub [41].
Figure 6: Experiments on large unit-disk graphs. Top: four graphs with 137 vertices each and whose hardness parameters span four orders of magnitude; the three hardest graphs are taken from [28]. In shades of blue, the probability $\braket{\hat{n}_{i}}$ to measure the $i$-th atom in the Rydberg state |$\ket{1}$ at the end of the $\mathcal{HP}$-linear protocol. Due to the high degeneracy of the MIS, some vertices are equally measured in the ground or excited state. Bottom: experimental results after post-processing to find an independent set of given size, with the MIS solution at the rightmost bin; error bars represent the 95% confidence intervals. The classical greedy algorithm (gray) fails to find the solution in all cases. The benchmark linear protocol (blue) reaches the MIS in the easiest case but fails for any harder instance. Finally, the $\mathcal{HP}$-linear protocol is able to find the MIS except for the hardest graph; the corresponding values of $P_{\text{MIS}}$ are indicated in the subtitles. Results for the two CD protocols (not shown) lie between those of the two linear protocols.
Supplemental Material
SI Ground state encoding of MIS
Given a simple undirected graph $G=(V,E)$ with vertex set $V$ and edges set $E$, a subset $S\subseteq V$ is called independent if no pair of vertices in $S$ are connected by an edge in $E$. In the maximum independent set (MIS) problem, we need to find the independent set of largest cardinality. Denoting by the binary variable $x_{i}$ the presence of the vertex $i=1,\ldots,N$ in $S$, it is proven that a MIS solution minimizes the cost function [42, 43]
$\mathcal{C}(x_{1},\ldots,x_{N})=-\sum_{i\in V}\Delta\,x_{i}+\sum_{ij\in E}V_{ij}x_{i}x_{j}$ (S1)
if $0<\Delta<V_{ij}$ for all $ij\in E$. In our setting, $G$ is a unit-disk graph that connects the nearest and next-nearest neighbors of a square lattice. Neglecting the interactions between farther atoms and for a non-planar graph (i.e., when crossing diagonals belong to it), the MIS condition becomes:
$0<\Delta_{\text{f}}<\frac{C_{6}}{(\sqrt{2}\,a)^{6}}=\frac{V_{0}}{8},$ (S2)
where $a$ is the lattice spacing (considered as a fixed parameter), $V_{0}=C_{6}/a^{6}$ is the interaction strength between adjacent atoms, and $\Delta_{\text{f}}=\Delta(t_{\text{f}})$ is the detuning at the final evolution time $t_{\text{f}}$. Obviously, for the Rydberg Hamiltonian to encode the MIS cost function at $t_{\text{f}}$, we also need $\Omega(t_{\text{f}})=0$, see Eq. (1).
In several investigations, it is implicitly assumed that the inequalities (S2) hold for all unit-disk graphs due to the sharp Rydberg blockade transition. However, in what follows we show that the condition $\Delta_{\text{f}}>0$ is necessary but not sufficient.
SI.1 Nonexistence of a lower bound on the detuning
The Van der Waals interaction decays quickly with the distance between atoms, but it cannot be neglected for distances close to the blockade radius. As an example, consider the graph that consists of three atoms in a row with a distance $2a$ between the extremities, see Fig. S1(a). In this case, the MIS solution is $(1,0,1)$, which corresponds to an energy $-2\Delta_{\text{f}}+V_{0}/64$. By listing the energy of all other configurations, we find the following conditions on $\Delta_{\text{f}}$ such that |$\ket{101}$ is the ground state of $H(t_{\text{f}})$:
$V0/64<\Delta_{\text{f}}<2\,V_{0}.$ (S3)
The upper bound is of course already satisfied in Eq. (S2) as it is a sufficient condition, but the value $V_{0}/64$ should be considered as a strict lower bound for any protocol; remark that this bound is mentioned in [20].
Figure S1: Examples of unit-disk graphs to study the lower bound on the final detuning. The MIS solution (filled circles) has a lower energy than the suboptimal configuration (orange circles) iff $\Delta_{\text{f}}>\Delta_{\text{LB}}$. (a) Reference graph used in Eq. (S3). (b) Graph with $N=3m+1$ vertices forming $m$ alternating upwards and downwards triangles. For any $\Delta_{\text{f}}>0$, there exists a value of $m$ such that the suboptimal configuration is energetically favorable. Finally, two non-planar graphs with very different lower bounds: (c) $\Delta_{\text{LB}}\approx V_{0}/1017$ ($\mathcal{HP}\approx 0.42$), and (d) $\Delta_{\text{LB}}\approx V_{0}/11.49$ ($\mathcal{HP}=15$).
The question is then to determine if there exists, for unit-disk graphs in a square lattice, a finite lower bound on the detuning to ensure the equivalence between the classical MIS solution and the ground state of the quantum Hamiltonian. Unfortunately, the answer is negative. In fact, consider the unit-disk graph in Fig. S1(b), and the following two sets of vertices:
  • $S_{0}$:
    the MIS solution with $m+1$ vertices on the main line (filled circles)
  • $S_{1}$:
    the $m$ vertices located at the tip of the alternating triangles (orange circles)
Taking into account all pairwise interactions, the corresponding energies read:
$E_{0}$ $=-(m+1)\,\Delta_{\text{f}}+\frac{V_{0}}{64}\sum_{i=1}^{m}\frac{m+1-i}{i^{6}},$ (S4a)
$E_{1}$ $=-m\,\Delta_{\text{f}}+\frac{V_{0}}{64}\sum_{i=1}^{\lfloor m/2\rfloor}\left(\frac{m+1-2i}{(4i^{2}-4i+2)^{3}}+\frac{m-2i}{(2i)^{6}}\right).$ (S4b)
For the ground state of the Hamiltonian to yield the MIS solution we must have $E_{0}<E_{1}$, i.e., the detuning has to satisfy the inequality
$\frac{V_{0}}{64}\left(\Sigma_{0}-\Sigma_{1}\right)<\Delta_{\text{f}},$ (S5)
where $\Sigma_{0,1}$ stands for the sums in the r.h.s. of Eq. (S4). However, one can check that the term $\Sigma_{0}-\Sigma_{1}$ is always positive and becomes arbitrarily large for increasing $m$. In particular, this difference is larger than 8 for $m\geq 9$, leading to a contradiction with Eq. (S2).
In conclusion, there does not exist a value of the final detuning such that the Rydberg Hamiltonian is ensured to encode the MIS solution of all unit-disk graphs.
SI.2 Statistical analysis of the bounds on the detuning
Figure S2: Lower bounds $\Delta_{\text{LB}}$ (purple) and upper bounds $\Delta_{\text{UB}}$ (yellow-green) on the detuning $\Delta_{\text{f}}$ such that the ground state of the Hamiltonian encodes the MIS solution. Calculations have been performed on 10 0001000010\,00010 000 non-isomorphic unit-disk graphs with 8 to 19 vertices. All but 4 graphs satisfy $\Delta_{\text{LB}}\leq V_{0}/12$ (dashed red line) while $\Delta_{\text{UB}}$ is always strictly larger than $V_{0}/8$. It is interesting to note that, in average, easy graphs exhibit much smaller lower bounds than hard graphs (red curve), see also Fig. S1(c-d). Since the average value of $\Delta_{\text{LB}}$ converges quickly with the hardness parameter, we are confident that these bounds can be effectively applied to graphs of any size.
As shown above, the final detuning has to be arbitrarily high for the Hamiltonian to encode in its ground state the MIS solution of some specific graphs. However, this is not the case in general, as most graphs obey well-defined conditions on $\Delta_{\text{f}}$. Based on an extensive analysis of 10 0001000010\,00010 000 graphs, see Fig. S2, we find that more than 99.9% of all graphs satisfy the following bounds:
$\Delta_{\text{f}}\in[\Delta_{\text{LB}},\Delta_{\text{UB}}]=\left[\frac{V_{0}}{12}\,,\frac{V_{0}}{8}\right].$ (S6)
Interestingly, this interval is very close to the best range of parameters found by a variational optimization, see [20, Fig. 6.1], with $\Delta_{\text{LB}}=28.91$ and $\Delta_{\text{UB}}=43.36$ MHz for QuEra’s Aquila device when $a=5$ µm. Note that a statistical analysis is also performed in [18, Sec. S2], where it is stated that about 99% of the MIS problems can be encoded in the ground state of a Rydberg Hamiltonian for well-chosen fixed experimental parameters.
In conclusion, we strongly recommend to consider the interval Eq. (S6) when setting the final detuning of any standard adiabatic protocol aiming to solve the MIS problem. A value which lies far outside this interval will likely lead to poor numerical or experimental results simply because the ground state of the Rydberg Hamiltonian does not correspond to the MIS solution.
SII Graph-dependent counterdiabatic driving
For convenience, we set $\hbar=1$ and $\phi(t)=0$ in what follows, but all results can be extended to the general scenario. In this case, the Rydberg Hamiltonian reads:
$H(t)=\frac{\Omega(t)}{2}\sum_{i}\hat{X}_{i}-\Delta(t)\sum_{i}\hat{n}_{i}+\sum_{i<j}V_{ij}\hat{n}_{i}\hat{n}_{j}\equiv\frac{\Omega(t)}{2}\ \hat{X}-\Delta(t)\ \hat{n}+\hat{V},$ (S7)
[…]
SII.1 Approximate adiabatic gauge potential
In order to suppress the transitions between eigenstates during the shorten adiabatic evolution, the gauge potential $\mathcal{A}$ in Eq. (5) must satisfy the following equation [26, Eq. (2)]:
$\left[\mathrm{i}\mkern 1.0mu\partial_{t}H-\left[\mathcal{A},H\right],H\right]=0.$ (S8)
Neglecting the interactions $\hat{V}$, it is shown in [27] that an exact solution of this equation exists:
$\mathcal{A}_{V\rightarrow 0}=-\frac{1}{2}\ \frac{\Omega\dot{\Delta}-\dot{\Omega}\Delta}{\Omega^{2}+\Delta^{2}}\ \hat{Y}.$ (S9)
This gauge potential leads to improved results compared to a simple adiabatic protocol, but one might expect even better results with a gauge potential that explicitly depends on the coefficients $V_{ij}$, hence being specific to the graph to be solved. In order to find such a generalized gauge potential, we consider the variational procedure that consists in minimizing the action [26, Eq. (8)]
$\mathcal{S}(\mathcal{A}_{\lambda})=\operatorname{Tr}\left(G^{2}(\mathcal{A}_{\lambda})\right)\quad\text{with}\quad G(\mathcal{A}_{\lambda})=\partial_{t}H+\mathrm{i}\mkern 1.0mu\ [\mathcal{A}_{\lambda},H].$ (S10)
Here, we use the ansatz $\mathcal{A}_{\lambda}=-\hat{Y}\lambda(t)/2$ since the one-body Pauli matrices $\hat{Y}_{i}$ are the only operators that contribute to the action while being experimentally implementable through the phase $\phi$ of the Rabi drive. One easily checks that $\mathcal{S}$ is quadratic in $\lambda$ and that the optimum value of this parameter is:
$\lambda=2\,\frac{\operatorname{Tr}\left(\partial_{t}H\cdot[\mathrm{i}\mkern 1.0mu\hat{Y},H]\right)}{\operatorname{Tr}\left([\mathrm{i}\mkern 1.0mu\hat{Y},H]^{2}\right)}.$ (S11)
Based on the algebraic properties of the Pauli matrices and of the trace operator, this expression can be developed and simplified as follows:
$\lambda=\frac{\Omega\dot{\Delta}-\dot{\Omega}\Delta+\dot{\Omega}\ T_{1}}{\Omega^{2}+\Delta^{2}-2\Delta\ T_{1}+T_{2}},$ (S12)
where the normalized traces $T_{1}$ and $T_{2}$ are given by:
$N\cdot T_{1}$ $=2^{-N}\operatorname{Tr}\left(\hat{X}\cdot[\mathrm{i}\mkern 1.0mu\hat{Y},\hat{V}]\right)=\sum_{i<j}V_{ij},$ (S13a)
$N\cdot T_{2}$ $=2^{-N}\operatorname{Tr}\left([\mathrm{i}\mkern 1.0mu\hat{Y},\hat{V}]^{2}\right)=\frac{1}{2}\sum_{i\leq j}\sum_{k\neq i,j}V_{ik}V_{kj}.$ (S13b)
As expected, the graph-dependent gauge potential $\mathcal{A}_{\lambda}$ is equal to $\mathcal{A}_{V\rightarrow 0}$ in the limiting case of vanishing interactions, with $T_{1}=T_{2}=0$. Due to the fast-decaying interaction between distant atoms, $T_{1}$ and $T_{2}$ do not scale with $N$ as they represent average values of (nearly) local functions. For instance, $T_{1}$ is exactly half the mean degree of the underlying graph if ones considers its adjacency matrix instead of the interaction terms in Eq. (S13a).
Finally, the driving amplitude and phase of the CD Hamiltonian as defined in Eq. (5) become:
$\Omega_{\text{CD}}=\sqrt{\Omega^{2}+\lambda^{2}},$ (S14a)
$\phi_{\text{CD}}=\arctan(\lambda/\Omega).$ (S14b)
Note that $\lambda$ and therefore $\dot{\Omega}$ should vanish at the beginning and at the end of the protocol to satisfy the boundary conditions $\Omega_{\text{CD}}(t_{\text{i}})=\Omega_{\text{CD}}(t_{\text{f}})=0$.
SII.2 Practical application of the counterdiabatic driving
Hardware implementation
It follows from Eq. (S14a) that $\Omega_{\text{CD}}$ is always larger than $\Omega$, so that it may exceed the maximum Rabi frequency $\Omega_{\max}$ of the quantum computer. In this case, we rescale the adiabatic function $\Omega$ by a factor $10<k<$ 1 to satisfy the hardware constraint at any time $t$ of the evolution:
$\max_{t}\left\{\Omega_{\text{CD}}(k\,\Omega(t))\right\}=\Omega_{\max}.$
Note that this is slightly different from rescaling $\Omega_{\text{CD}}$ directly, as the relations (S12, S14) would not hold in that case.
Variational parameter
Since Eq. (S12) is a generalization of the counterdiabatic term proposed in [27] through the addition of $T_{1}$ and $T_{2}$, it is interesting in an optimization context to consider a continuous interpolation between these two protocols. One possibility is to consider the interaction strength $V_{0}$ as a free parameter in the range $[0,C_{6}/a^{6}]$ when calculating the CD driving. To avoid any confusion with the fixed interactions in the main Hamiltonian, we denote by $\nu$ this free parameter. Numerically, this is equivalent to replace $T_{1}$ and $T_{2}$ in Eq. (S12) by the following expressions:
$T_{1}$ $=\nu\,T_{1}^{(0)},$ (S15a)
$T_{2}$ $=\nu^{2}\,T_{2}^{(0)},$ (S15b)
where $T_{1}^{(0)}$ and $T_{2}^{(0)}$ are the traces evaluated at $V_{0}=1$ in Eq. (S13).
SIII Graph datasets and numerical implementation
In the following sections, we describe two sets of unit-disk graphs $G=(V,E)$ that are used in this study; data are available on GitHub [41]. In all examples, we consider an underlying square lattice where the distance between nearest-neighbors is $a=1$ and where any pair of vertices in $V$ is connected by an edge in $E$ if the corresponding distance is smaller than or equal to $\sqrt{2}$. Then, we detail the numerical method that is used to simulate the evolution of the quantum state under the action of the Rydberg Hamiltonian. Finally, we provide the optimized schedule functions for the linear and for the counterdiabatic protocol with 6 and 3 free variables, respectively.
SIII.1 Graph datasets
Small graphs
From 10 0001000010\,00010 000 non-isomorphic graphs randomly generated, we choose 500 graphs such that:
  • there are exactly 50 graphs for each order from 8 to 17
  • the hardness parameters are distributed as evenly as possible in a log-scale from 0.375 to 12; in particular, no two graphs share the same hardness parameter, and exactly 100 graphs are contained in each bin $[3\cdot 2^{i-4},3\cdot 2^{i-3})$ with $i=1,\ldots,5$
  • the distribution of the graph orders is as uniform as possible within each such bin
The result is shown in Fig. S3. The main idea behind these rules is to generate a set of graphs for which the correlation between the hardness parameter and the order is as small as possible. In fact, in this study we want to show that the optimum parameters for the adiabatic evolution depend strongly on the hardness of the graphs and only weakly on their order.
Figure S3: Visualization of the hardness parameters $\mathcal{HP}$ and the orders of the 500 small graphs. Due to the combinatorial definition of the hardness parameter, graphs with fewer vertices (down purple triangles) cannot generate very hard MIS instances. This explains why graphs of order 13 and more (up yellow-green triangles) are overrepresented in the range $\mathcal{HP}\in[6,12]$.
Toy graphs
For completeness, we display in Fig. S4 the 11 “toy graphs” as defined in [28, Appendix C]:
  • the underlying square lattice has $4\times 3$ sites
  • the order of the graphs is 9 or 10
  • there is only one MIS solution
Figure S4: The 11 “toy graphs” sorted by the number of vertices (9 or 10) and by hardness parameter (between 2 and 3). The unique MIS is highlighted with dark blue circles. To refer to a specific graph, we generate the three-nibble identifiers as follows: the $i$-th nibble (four-bit aggregation) is the hexadecimal symbol that represents the absence or presence of a vertex in the 4 sites of the $i$-th row (from top to bottom).
SIII.2 Numerical implementation
Several software development kits (SDK) exist to simulate the Rydberg evolution: Bloqade [44] and Amazon Braket [31] for QuEra’s hardware, or Pulser [32] based on QuTiP for Pasqal’s technology, just to name a few. Nevertheless, we developed our own routines to get a full control over all parameters and, more importantly, to fit exactly our needs. In fact, optimizing hundreds of graphs that consist of $8\leq N\leq 17$ vertices for several protocols requires intense (classical) computing power, so that we have to find the perfect balance between accuracy and speed. For example, for each graph of the datasets we precompute all time- and protocol-independent terms in the Hamiltonian, namely the $\hat{X}$, $\hat{Y}$, $\hat{n}$ and $\hat{V}$ operators, see Eq. (S7). We also use an explicit Runge-Kutta method of order 2 [45] to integrate the time-dependent Schrödinger equation, which turns out to be sufficiently accurate and much faster than the default order 8 method (DOP853 in Python or Vern8 in Julia).
Figure S5: Numerical simulations of the Rydberg Hamiltonian for the toy graph “7DE” and the ACQC protocol defined in [27] with $t_{\text{f}}=1$ µs. The height of the bars represents the probability to find the atom $i$ in the state |$\ket{1}$ at the end of the evolution, with $i=1,\ldots,9$ from top to bottom and left to right in the graph. The MIS solution (1, 3, 7, 9) is clearly visible in this bar chart. We verify that our implementation (light colors) is perfectly equivalent to that of Bloqade (dark colors); note that, depending on the SDK, the phase has to be multiplied by minus one in Eq. (S14b). We also observe that the simulations in the full Hilbert space (blue bars) match those performed in the blockade subspace of nearest neighbors (orange bars). However, the unit-disk subspace approximation $R_{\text{s}}=R_{\text{b}}$ (green bars) is clearly in favor of the MIS solution and thus cannot be considered as satisfactory. This is even more noticeable when comparing the probability to reach the ground state at the end of the evolution for each blockade subspace: 0.49(4), 0.49(6), and 0.70(0) respectively.
Blockade subspace
Besides the above-mentioned implementation aspects, a crucial ingredient is the correct use of the blockade subspace: when the interaction energy between Rydberg states is much larger than the Rabi strength, the exact dynamics in the full Hilbert space is well approximated by the dynamics in the subspace where only one Rydberg excitation is allowed between nearby atoms [46]. In numerical simulations, the states |$\ket{11}_{ij}$ are therefore neglected if the distance between $i$ and $j$ is smaller than a given subspace radius $R_{\text{s}}$, which results in much faster computations. The default setting in most programs is $R_{\text{s}}=R_{\text{b}}$, so that the quantum evolution is restricted to the independent sets in unit-disk graphs. However, this approximation is too coarse to get accurate results, see Fig. S5: neglecting the states |$\ket{11}$ for next-nearest neighbors (extremities of a diagonal) leads to success probabilities that are overestimated by up to 50%. Therefore, it is crucial to simulate either the full Hilbert space or the blockade subspace where only nearest neighbor Rydberg states are neglected.
Optimization algorithm
In this study, we focus on the probability to reach a MIS at the end of the adiabatic evolution and not on the normalized mean energy (approximation ratio) of the final state, as it often the case for this problem. In fact, all results here revolve around the hardness parameter, which is intrinsically related to the two lowest-energy levels and not to the full energy spectrum. We therefore define the score function to be maximized as the overlap between the quantum state |$\ket{\psi}$ at the end of the evolution and all MIS solutions of the graph under investigation:
$P_{\text{MIS}}(\Theta)=\sum_{i=1}^{D_{|\text{MIS}|}}\big{|}\,\braket{\psi(\Theta,t_{\text{f}})\,}{\,e_{i}}\,\big{|}^{2},$ (S16)
where $e_{i}$ are the vectors of the canonical basis that correspond to the $D_{|\text{MIS}|}$ configurations of the (possibly) degenerate solution. The parameters $\Theta$ to be optimized depend on the protocol, e.g., $\Theta=(\tau_{\text{i}},\tau_{\text{f}},\Delta_{\text{i}},\Delta_{\text{f}})$ for the simple linear schedules. Since the evolution is deterministic, any standard optimization algorithm may be applied. We choose the simplex search method [47], which turns out to be robust and very efficient in our setting. We set $t_{\text{f}}=1$ µs, which leads to success probabilities that are well distributed over the whole unit interval for the 500 graphs of the dataset.
SIII.3 Optimized adiabatic schedules
The optimum parameters $\Theta$ depend principally on the hardness parameter of the graphs—and only marginally on their order $N$ for instance—so that an effective model $\hat{\Theta}(\mathcal{HP})$ can be fitted to the 500 small graphs. The simplest linear schedules are illustrated in Fig. 2 in the main text. In Fig. S6 and Fig. S7, we display the optimized schedules for a more advanced linear protocol with 6 free parameters and for the generalized counterdiabatic protocol with 3 free parameters. Finally, we benchmark these hardness-dependent protocols against those described in [27]: the success probability is increased for each of the 500 graphs of the dataset, see Fig. S8.
Figure S6: Optimized piecewise linear schedules. The schedules depend on the hardness of the graph and are illustrated here for $\mathcal{HP}\in\{0.5,1,2,4,8\}$ (thin lines, from green to maroon) and in the limit $\mathcal{HP}\to\infty$ (dark maroon thick line). In this latter case, the optimized parameters are $\tau_{\text{i}}\approx 0.22\,t_{\text{f}}$ and $\tau_{\text{f}}\approx 0.12\,t_{\text{f}}$ for the Rabi amplitude (left) and $\tau_{\text{m}}\approx 0.34\,t_{\text{f}}$, $\Delta_{\text{i}}\approx-0.83\,\Delta_{\text{UB}}$, and $\Delta_{\text{f}}=\Delta_{\text{UB}}$ for the detuning (right).
Figure S7: Optimized generalized counterdiabatic schedules. In this case, the functions $\Omega$ and $\phi$ depend not only on the hardness parameter but also on the duration of the protocol (here, $t_{\text{f}}=1$ µs) and on the value of the traces $T_{1}$ and $T_{2}$, see Eq. (S12) and Eq. (S14). The thin lines are based on the average values of the 500 small graphs: $\overline{T}_{1}\approx 0.93$ and $\overline{T}_{2}\approx 1.46$. The light green area represents the maximum variation of the phase for a fixed hardness parameter $\mathcal{HP}=0.5$.
Figure S8: Probability of finding a MIS in the 500 small graphs for the main protocols. Compared to the benchmark protocols, $P_{\text{MIS}}$ is improved in average by a factor of 3.6 for the piecewise linear schedules (left) and by a factor of 1.5 for the couterdiabatic schedules (center) developed in this study. Except for rare instances, the $\mathcal{HP}$-linear protocol always beats the $\mathcal{HP}$-ACQC protocol (right). Note that $P_{\text{MIS}}$ depends strongly on $\mathcal{HP}$ but not on $N$, with low values for the hard graphs (maroon) and nearly perfect results for the easy ones (green).